OpenStudy (anonymous):

A corner store sells two kinds of baked goods: cakes and pies. A cake costs $5 and a pie costs $7. In one day, the store sold 15 baked goods for a total of $91. How many cakes did they sell? A. 7 cakes B. 4 cakes C. 8 cakes D. 5 cakes

OpenStudy (anonymous):

i know it is not (C)

OpenStudy (anonymous):

i think (A)

OpenStudy (anonymous):

c+p=15 5c+7p=91

OpenStudy (anonymous):

so it is (A) right

OpenStudy (anonymous):

@eliassaab @hartnn

hartnn (hartnn):

could u solve those two equations??

OpenStudy (anonymous):

no

hartnn (hartnn):

multiply 5 to 1st equation and subtract it from 2nd,then tell me what u get.

OpenStudy (anonymous):

X what ?

hartnn (hartnn):

multiply 5 to 1st equation to get 5c+5p=75 now subtract this from 2nd equation

OpenStudy (anonymous):

16

hartnn (hartnn):

thats 2p=16 so p=8 now can u find c from c+p=15?

OpenStudy (anonymous):

7

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

(A)